### v2

# FilterPrompt: Guiding Image Transfer in Diffusion Models

Abstract
In controllable generation tasks, flexibly manipulating the generated images to attain a desired appearance or structure based on a single input image cue remains a critical and longstanding challenge. Achieving this requires the effective decoupling of key attributes within the input image data, aiming to get representations accurately. Previous research has predominantly concentrated on disentangling image attributes within feature space. However, the complex distribution present in real-world data often makes the application of such decoupling algorithms to other datasets challenging. Moreover, the granularity of control over feature encoding frequently fails to meet specific task requirements. Upon scrutinizing the characteristics of various generative models, we have observed that the input sensitivity and dynamic evolution properties of the diffusion model can be effectively fused with the explicit decomposition operation in pixel space. This integration enables the image processing operations performed in pixel space for a specific feature distribution of the input image, and can achieve the desired control effect in the generated results. Therefore, we propose FilterPrompt, an approach to enhance the model control effect. It can be universally applied to any diffusion model, allowing users to adjust the representation of specific image features in accordance with task requirements, thereby facilitating more precise and controllable generation outcomes. In particular, our designed experiments demonstrate that the FilterPrompt optimizes feature correlation, mitigates content conflicts during the generation process, and enhances the model’s control capability.
Keywords:
Visual Prompt, Image Transfer, Image-to-Image Translation, Explicit Decomposition, Controllable Generation, Diffusion Models
Figure 1: Comparison of image transfer tasks. Our approach FilterPrompt enables image transfer in multiple domains at local, object-centric, and full-graph levels. Compared to other works [1, 50, 52], our approach can help the model better preserve the geometric properties of structural images while maintaining consistent color distribution and texture features with appearance images.
1 Introduction
In the classic realm of Controllable Image Generation, achieving flexible control over the characteristics, attributes, or style of generated images remains a focal point of research. Researchers typically focus on acquiring low-dimensional feature representations of high-dimensional image features while minimizing the issues of information loss and attribute entanglement [5]. Alternatively, they seek solutions that offer greater interpretability in the generation process, promoting a more stable and efficient model training process, thus fostering further advancements in the field of image generation.
Controllable generation typically follows two approaches: 1. Some works involve disentangling the characteristics of an input image in the feature space and obtaining feature representations relevant to the control objective. Subsequently, the network regulates the degree to which these feature representations are expressed in the generated image, employing a diverse array of meticulously designed loss metrics [8, 4, 51, 29, 46, 16]. 2. Another category of work involves incorporating a conditioning mechanism into the architecture of the model. These works improve the capacity of the model to integrate control conditions while learning the target style domain’s data distribution. Consequently, the generated image exhibits a style map akin to the training set, yielding an artistic effect that aligns with the input content image [34, 21, 30, 43].
The above controllable generation ideas have their limitations. On the one hand, there will be an unavoidable loss when mapping different data domains to the same feature space. Simultaneously, it is impossible to ensure that there is no entanglement of attributes among the representations [48, 6]. On the other hand, using machine learning algorithms to train style mappings may not have the same level of interpretability as traditional mathematical modeling approaches, and the training of such models often requires expensive datasets [7, 45].
Hence, unlike previous endeavors focused on refining algorithms and models within the feature space, we redirect our focus toward the pixel space. Intuitively, certain semantic features, often indistinguishable in models, exhibit discernible distribution discrepancies visible to the naked eye in the pixel space. Moreover, manipulating abstract feature representations within the feature space may inadvertently impact the entire result while executing image processing operations on input data within the pixel space is more intuitive and controllable.
Figure 2: Enhance network performance with filtering operations. Filtering operations can enhance image quality. When neural networks extract specific image features, strategically incorporating filtering operations upstream enables targeted suppression or enhancement of particular feature distributions.
Following experimental comparisons of various mainstream generative models, we observe that the diffusion model possesses properties of input sensitivity and dynamic evolution which are very sort for image processing operations in the pixel space. This integration enables image processing operations performed in pixel space for a specific feature distribution of the input image and can achieve the desired control effect in the generated results. Therefore, we name this approach FilterPrompt.
To validate the aforementioned assertion, we build a task framework for experimental testing based on the existing pre-trained model and use it to demonstrate the impact of FilterPrompt on the control effect of appearance features and structure features of the generated image. Finally, we conduct quantitative analyses of the generated results. These analyses demonstrate that the FilterPrompt optimizes feature correlation, mitigates content conflicts during the generation process, and enhances the model’s control capability. We believe that our approach can be easily generalized to more complex combinations. At the same time, FilterPrompt has the characteristics of requiring no additional training and being highly interpretable.
In summary, our contributions are shown as follows: 1. We introduce a new approach, FilterPrompt, aimed at optimizing model control effects. Our approach can be customized for various input data based on specific task requirements and combined with the diffusion model to achieve the expected control effect. 2. We analyze how FilterPrompt facilitates the desired control effect within the diffusion model framework. Additionally, we have designed experiments to provide explanations to demonstrate the feasibility of our approach. 3. Our experiments encompass a range of tasks, including local, object-centered, and full-graph appearance transfer, as well as image style transfer. We present the application of FilterPrompt in these tasks and compare its performance with various other models. Through experimentation, we substantiate the efficacy of our approach in image transfer.
2 Related Work
2.1 Controllable Image Generation in Diffusion Modeling
In the development of the diffusion model, the research on the Controllable Image Generation task has roughly three stages. The first stage of the work is mainly based on the iterative denoising process and achieves controllable generation effects by rationally using the input image to generate a deterministic guided generation paradigm [9, 33, 31]. This stage of the work controls the semantic similarity between the generated image and the input image by influencing the proportion of information mixing in the denoising network of U-Net. The second phase of the work is based on the image generation strategy guided by the display classifier. Optimization is performed by adding gradient information from the classifier to the loss function. This idea first originated from the concept of "Classifier Guidance," proposed by[42] and further advanced in[11, 37]. Since then, numerous studies have broadened the scope of classifiers, extending the classification guidance of the diffusion model to encompass diverse modalities such as text, images, and other multi-modal data [2, 22, 47, 16, 28]. The third stage marks the era of large models based on implicit classifiers. To address the issue of declining diversity in Classifier Guidance, the "Classifier-Free Guidance" strategy [18] emerges later. This approach involves decomposing the gradient guidance from the explicit classifier into two components. One component is an unconditionally generated gradient prediction model, akin to the conventional DDPM [17]. The other is a gradient prediction model based on conditional generation, conceptualized as a U-Net network with an overlay of a cross-attention mechanism. The success of this approach has catalyzed the evolution of a variety of subsequent image editing technologies. These technologies utilize the diffusion model as a foundational framework and integrate the attention mechanism, resulting in notable progress in the application of the diffusion model across diverse fields. Notable projects in this domain include DALLE-2 [36], DreamFusion [35], Stable Diffusion Model (SDM) [38], and more.
Here, our primary focus lies on Grounded Generation and Layout-driven Generation within the context of the controllable generation problem. Representative works in this area include GLIGEN [26], ControlNet [52], and IP-Adapter [50]. We aim to investigate the generative capabilities of diffusion models in addressing semantic-level conditional guidance, particularly in scenarios with limited sample sizes.
2.2 Explicit Decomposition
Explicit decomposition is aimed at breaking down the representation of data or a model into simpler, more independent components or factors. Specifically, this process involves splitting a high-dimensional representation space into multiple low-dimensional subspaces, each responsible for encoding a specific aspect or attribute of the data. Through explicit decomposition, neural networks can more easily understand various aspects of the data, such as geometry, color distribution, texture, etc., in the image [19].
Traditionally, filtering algorithms have been considered the explicit decomposition approach as they can break down input data into components at different frequencies or spatial scales. Examples of common filters include the Gaussian filter, Sobel filter [41], Adaptive filter [20], and Gabor filter [12, 39]. These filters can weigh data at different frequencies or scales to suppress or enhance specific features, as illustrated in Figure 2. Performing preprocessing operations helps neural networks better understand the structure and features of data in order to acquire more refined representations.
3 Prerequisite
3.1 Prompt Impact on the Iterative Process of Diffusion Models
Specifically, the forward diffusion defines a known Gaussian translation process. So the image intermediate quantity $x_{t}$ at each moment of the forward process can establish a unique relationship with the input image $x_{0}$ as:
$q(x_{t}|x_{0}):=N(x_{t};\sqrt{\bar{\alpha_{t}}}x_{0},(1-\bar{\alpha_{t}})I),\;\;\;x_{t}=\sqrt{\bar{\alpha_{t}}}x_{0}+(1-\bar{\alpha_{t}})\epsilon$ (1)
The reverse denoising process of the diffusion model can be seen as a migration process to the target data distribution. During this process, the model will continuously try to reduce noise and be guided by the condition $c$ to restore the structure or characteristics of the specified data. Every migration process can be expressed as:
$x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}\left(x_{t},c,t\right)\right)+\sigma_{t}z$ (2)
The focus of this process is the network’s prediction of noise distribution: $\epsilon_{\theta}\left(x_{t},c,t\right)$. This prediction is affected by the current moment $x_{t}$ and condition $c$, and condition $c$ comes from the external reference image $y$ mapping result. Therefore, if we perform filters $f_{\gamma}$ on either side of the input like making $c=\mathrm{Encoder}(y)$ become ${c{{}^{\prime}}=~{}\mathrm{Encoder}(f_{\gamma}(y))~{}}$, it will affect the prediction results of the noise distribution $\epsilon_{\theta}$, and even affect the migration direction of the generated distribution at that time node as illustrated in Figure 3. The impact caused by conditions $c$ or the input content image $x_{0}$ would be reiterated at each sampling instance, and the minor changes introduced by filters would be involved in the entire generation process.
In the subsequent experiment depicted in Figure 8, we observe that, after applying filtering operations to a specific feature distribution in pixel space, the degree of expression for that feature distribution aligns with expectations. Moreover, the filtering operations targeting a specific feature distribution does not affect the expression of other features. This indicates that the influence of filters is independent in the feature space from the encoding of other representations. It can be seen that prompts in pixel space offer a lightweight and convenient way to optimize the entanglement between feature representations in the diffusion model.
3.2 Static Generation vs. Dynamic Generation
We pay attention to the fact that in the sampling inference stage, the diffusion model has different dynamic evolution properties from the traditional generative model. In the basic theory of the diffusion model, the process of generating an image starts from a simple noise image [17], and through multiple iterations, the noise in the image is gradually removed until a final clear image is generated. From this perspective, the diffusion model generation process is continuous in time, and the image is generated through a gradually changing process. Therefore, we believe that the gradual evolution of the diffusion model makes it a dynamic generative model.
Figure 3: Part (a) shows the illustration of filter’s impact on the corresponding sampling inference stages in the static generative model and the dynamic generative model. Part (b) gives a comparison of the results obtained by applying filter to some works of generative models. These include FastStylePredict [15], which is a fast style transfer algorithm based on CNN, CycleGAN [55], DualStyleGAN [49] based on GAN and AE architecture, InST [53] and IP-Adapter [50] based on SDM architecture and the effects of our work. Comparing the last column of figures, we can intuitively see that filter operation has a more significant impact on diffusion models.
In the early days, the fast style transfer algorithm [14, 25, 15] is based on convolutional neural network (CNN), AutoEncoder (AE), and Generative Adversarial Network (GAN) [10, 55, 49], and their fitting effect on the target data distribution depends on the design of the training process. During the training process, the model attempts to capture the statistical characteristics of the entire data distribution and obtain the maximum likelihood representation of the data distribution. When training is completed, the generated samples exist statically in the latent space. Therefore, we consider this type of model to be the static generative model.
We noticed that prompts in pixel space may exhibit certain limitations in previous static generation models, but it can be well combined with dynamic generation models. In the style transfer task shown in Figure 3, We applied the same level of blur interference to the input data of various generative models. The results indicate that the impact of blur prompt on static generative models is relatively weak, while the outcomes of diffusion models show significant changes. We believe that for static generation models, using prompt on input data during the sampling inference stage will only have a slight impact on the mapping position of this input in the static distribution. And this position is close to the mapping result of the original image $x$. Therefore, prompts in pixel space will not significantly affect the generation results of static generation models. However, for diffusion models, as indicated by Equation 1, there is a predefined long-term dependency between the current time node image $x_{t}$ and the input image $x_{0}$, in the forward process. Therefore, the impact of using prompts on the input content image $x_{0}$ will be executed again at each sampling, and the minor changes caused by prompts will participate in the entire generation process, thereby exerting a more significant influence on the generation results.
4 Proposed Approach
It is noted that distinguishing and processing analogous feature representations in feature space leads to significant computational overhead. Consequently, we propose the adoption of our pixel-space methodology, termed FilterPrompt. This approach directly manipulates the frequency or distribution characteristics of specific image attributes, thereby influencing the subsequent expression levels of the representation. FilterPrompt offers an intuitive and straightforward approach. Fundamentally, within the diffusion model framework, our approach facilitates deterministic guidance on the generation process through the integration of filtering operations.
To provide a more detailed exposition and validation of the aforementioned statement, the arrangement of this chapter is as follows: In Section 4.1 We provided a general description of the FilterPrompt. In Section 4.2 To verify and discuss the impact of our approach on the control effect of diffusion models, we provide the experimental setup, including the construction of data sets and the framework of baseline. In Section 4.3 we give the effects of various FilterPrompts on the baseline. In Section 4.4 we show the visual explanation of FilterPrompt on the results of each sampling in the denoising generation process.
4.1 FilterPrompt
Given the dynamic and evolving nature of the diffusion model, an intuitive approach emerges: utilizing prompts on the input images to modulate the expression levels of specific features. We term this approach FilterPrompt.
Specifically, we leverage IP-Adapter[50] to obtain the conditional encoding $C_{s}$, which governs the rendering attributes of the generated result from the reference image $x_{style}$. Simultaneously, we employ ControlNet[52] to obtain the conditional encoding $C_{c}$, responsible for regulating the geometric attributes from the reference image $x_{content}$. $C_{s}$ and $C_{c}$ can be regarded as a collection of features from reference images.
$C_{s}=\left\{feature_{\{m\}(x_{\{style\}})}\right\}_{\{m=1\}}^{\{M\}},C_{c}=\left\{feature_{\{n\}(x_{\{content\}})}\right\}_{\{n=1\}}^{\{N\}}$ (3)
However, due to the inability to guarantee disentanglement between the features of $C_{s}$ and $C_{c}$, directly using them for guided generation in the latent space would inevitably result in content conflicts. Drawing upon the successful performance of prompts demonstrated in the dynamic generation model above, we try to alleviate these conflicts by leveraging FilterPrompt. Moreover, features often exhibit distinct differences in frequency or distribution in pixel space. This fact empowers us to tailor the design of FilterPrompt to meet specific task prerequisites. For example, integrating a color-removing prompt $F$ will nullify color features, leading to the subsequent alterations in coding:
$C_{s}^{\prime}=\left\{feature_{\{m\}(F(x_{\{style\}}))}\right\}_{\{m=1\}}^{(M-k)}\quad,C_{c}^{\prime}=\left\{feature_{\{n\}(F(x_{(content)}))}\right\}_{\{n=1\}}^{\{N-k\}}$ (4)
At the same time, the pixel range affected by $F$ is intuitive and convenient, so we can estimate the impact of it on the migration of Gaussian distribution in Equation 2 and change the iterative process into:
$x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}\left(x_{t},c_{\{s\}}^{\prime},c_{\{c\}}^{\prime},t\right)\right)+\sigma_{t}z$ (5)
In conclusion, we advocate for the adoption of our innovative approach, FilterPrompt, which directly manipulates the frequency or distribution characteristics of specific image attributes, thereby influencing the subsequent expression levels of the representation. This approach helps the diffusion model demonstrate more powerful control capabilities when dealing with complex generation tasks.
4.2 Architecture Description
Framework for Experiments: In order to verify and discuss the impact of our approach on the control effect of diffusion models, our experiment are built based on combined filtering operations, ControlNet [52] and IP-Adapter [50], as shown in Figure 4. This means we combine ControlNet with IP-Adapter as the baseline. Specifically, inspired by the concept of decoupled cross-attention in IP-Adapter, we apply a similar methodology. Subsequently, we map the low-level features in the style image to a global embedding $C_{s}$ and concatenate it with the default text prompt embedding $C_{t}$ of SDM. This process can be described as $X_{t}=C_{t}\bigoplus C_{s}$. These two parts in hidden state $x_{t}$ are processed separately at each denoising generation. A portion of $x_{t}$ is delegated to ControlNet, which projects the latent distribution into a fused distribution controlled by high-level features $C_{c}$. The global embedding of another part in $x_{t}$ utilizes the IP-Adapter for decoding, unfolding, and guiding the generation of low-level features. We use $x_{t-1}$ to represent hidden state predicted at the next moment in Equation 2. The intermediate hidden states obtained from both processes are weighted and summed according to Equation 6, achieving the effect of unifying representations related to Content and Style into the latent space of SDM as shown in Equation 6.
$X_{t-1}=\alpha\cdot ControlNet_{hiddenStates}+\beta\cdot\lambda\cdot IP_{hiddenStates}$ (6)
where $\alpha$, $\beta$ are weight control factors and $\lambda$ is scale control factor.
In experiments 4-2-1 and 4-2-2, $\alpha=1$, $\beta=1$, $\lambda=7.5$, and the random seed of the diffusion model is 42. In experiments 4-2-3, 4-3, and quantitative analysis, $\alpha=1.2$, $\beta=0.6$, $\lambda=7.5$, and the random seed of the diffusion model is 52.
Figure 4: Framework for experiments. The experiment uses ControlNet and IP-Adapter as the baseline and adds combined filtering operations as the expansion. We mapped low-level features in style images to global embeddings as $Cs$, concatenating them with SDM default text prompt embeddings $Ct$. The denoising generation processes these parts separately. A segment is managed by ControlNet, projecting latent distributions into a fused distribution controlled by high-level features that is $Cc$. The other part uses IP-Adapter for decoding and guiding low-level feature generation. Intermediate hidden state $x_{t-1}$ from both processes are weighted and summed every sampling time.
4.3 Effects of Various FilterPrompts on Baseline
For experiments, we define structure to represent the geometric features in the content image and appearance to denote the rendering features derived from the style image’s color distribution and pixel-level texture. Next, we assess and compare the performance variations of the baseline on structural features from content images and appearance features from style images before and after applying FilterPrompt (see Figure 5).
Figure 5: Comparison of the effect of existing approaches and FilterPrompts on structural control. The generated results demonstrate that the combined filters in FilterPrompt(4) have the best preservation effect of structural information. Specifically, these filters enable high-fidelity reproduction of critical vehicular details such as the exhaust window and headlights, which existing preprocessing methodologies fail to replicate.
Firstly, we perform our approach on the ControlNet path, which itself is used for control structure information in the baseline. The comparison of results among existing image preprocessing approaches and FilterPrompt indicates that our approach retains more details in the generated results and brings the colors closer to the ideal appearance. Meanwhile, comparisons between text control and FilterPrompt indicate that the impact of our approach is more pronounced, as shown in Figure 5. In the various FilterPrompt configurations shown in the figure, the combinations (using the ITV-R 601-2 luma transform method for decolorization, inversion, and contrast enhancement of the grayscale image) in Figure 5-FilterPrompt(4) demonstrates the best preservation effect of structural information. The purpose of these filters is to change the brightness distribution of the picture to make the outline structure clearer. Therefore, FilterPrompt exhibits strong intuitiveness and interpretability.
Figure 6: Effect of different FilterPrompts on appearance image. Among the results, the Sharpen filter enhances the expression of fine strokes, while the Gaussian filter blurs detailed stroke information. This illustrates that FilterPrompt can have a discernible impact on appearance information, aligning well with our expectations.
Figure 7: Impact of different kernel sizes in FilterPrompt on the generated results. In this example, we utilize FilterPrompt(4) mentioned before on the structure image. Simultaneously, a Gaussian filter is applied to the appearance image. The outcomes highlight the effectiveness of FilterPrompt(4) on the structure image in preserving the geometric attribute information of the bronze. Additionally, increasing the Gaussian kernel size helps reduce the representation of redundant pattern information in the appearance image, thus addressing content conflicts in the generated results.
Additionally, we also conduct a comparative analysis of the IP-adapter path, which controls appearance information, both before and after the implementation of FilterPrompt. According to the results in Figure 6, we observed that after applying FilterPrompt designed for noise processing to the current appearance image, it significantly affects the detailed stroke information of the generated image. At the same time, the generated effects are consistent with the characteristics of the applied filters themselves. For example, the Sharpen filter enhances the expression of fine strokes, while the Gaussian filter blurs detailed stroke information. This experimental result verifies the capability of FilterPrompt to produce the expected control effects on appearance.
Taking a step further, we explore the control effect of FilterPrompt on both paths in the baseline. In the task that requires high protection of geometric features (Figure 7), the aim is to transfer a bronze sketch to a photo with a specified appearance image without altering geometric features from the structure image. We use the same prompts in Figure 5-FilterPrompt(4) for ControlNet path, to enhance control over structure features. However, the initial results reveal redundant pattern semantic expressions on the bronze surface. So, we apply Gaussian filter for the IP-Adapter path to help suppress high-frequency information. It can be seen from the histogram in Figure 7 that Gaussian filter hardly changes the color distribution but will destroy the continuous distribution of high-frequency information. The generation results show that when the Gaussian kernel increases to 4, the redundant pattern features in the generation results significantly fade.
4.4 Explanation
Figure 8: Sampling inference process. After applying our approach FilterPrompt, the underlying texture in the sampled images changes from a distribution resembling arc patterns to a point-like distribution. Additionally, as shown in the enlarged illustration on the right, it is evident that the use of filters consistently disrupts the expression of redundant pattern features.
We previously discussed that FilterPrompt, when viewed from the perspective of dynamic generation, serves as a guiding force in steering the Gaussian distribution of current data toward the target distribution during migration and diffusion. This controlled influence on distribution migration is elucidated in this subsection by examining the effects of FilterPrompt at various stages during the sampled inference process.
As shown in Figure 8, we perform a detailed analysis of the sampling inference stage for the task of converting a bronze sketch to a photo in the preceding section. Notably, the utilization of Gaussian FilterPrompt interrupts the expression of redundant texture features. The sampled results vividly illustrate that detailed representations of arc patterns initially present in the early stages are weakened by Gaussian FilterPrompt, manifesting as point-like distributions. This observation suggests that Gaussian FilterPrompt disrupts the continuous semantic expression of patterns. Concurrently, subsequent texture generation doesn’t emerge the negative impact of full-graph blurring, so it illustrates the property that FilterPrompt is only effective for specific feature distributions.
The third row of the evolution sequence in Figure 8 illustrates the impact of applying Gaussian FilterPrompt only in the first 12 steps, where we observe the absence of redundant pattern features in the final generation result. A set of comparisons in lines 4 and 5 showcase the effectiveness of Gaussian FilterPrompt. In the first 6 steps, only Controlnrt is utilized to regulate the structural layout. Introducing IP-adapter in the 7th step guides appearance features. The final results reveal that, even if Gaussian FilterPrompt does not initially suppress redundant features, it remains effective in later stages during the generation of detailed textures. The above findings highlight that the impact of filtering on the diffusion process is intuitive, controllable, and predictable.
5 Results
In this part, we qualitatively and quantitatively analyze the control effect impact of our approach by using FilterPrompt(4) in Figure 5. The generated results demonstrate our approach can be conveniently generalized to various data domains and controllable generation tasks.
Figure 9: Comparison with other works.
Table 1: Metrics evaluation. The results demonstrate that FilterPrompt achieves better performance in preserving structure, shape, and edge similarity, as well as in maintaining feature distribution similarity, texture differences, image quality, and color histogram correlation. We highlight the best value in red, and thesecond-best value in yellow.
Structure Preservation Shape and Edge Feature Distribution Texture Quality Color Correlation
SP↑ CD↓ FID↓ GLCM↓ PSNR↑ CHC↑
Cross-Image 0.7791 5.4133 245.0973 0.1376 9.4278 0.9357
IP-Adapter 0.8313 4.0967 210.2189 0.1619 9.5546 0.8004
Baseline 0.8547 3.3027 222.8576 0.1618 10.5011 0.9364
Ours 0.8799 2.8092 215.8267 0.1072 10.5594 0.9405
5.1 Quantitative analysis
The essence of the appearance transfer task lies in effectively blending the high-level features found in structural images with the low-level features present in appearance images. High-level features in structural images encompass their geometric properties and semantic characteristics, whereas low-level features in appearance images encompass aspects such as color distribution and texture features. Our quantitative analysis focuses on six distinct types of appearance transfer tasks: cat to cat, cat to dog, cat to wild, bird to bird, airplane to bird, and car to car. It’s essential to note that all structure-appearance image pairs utilized in these experiments share the same style: Photographs.
By referring to some works on image transfer tasks [40, 27, 13, 1], we selected six metrics that align with the objectives of our quantitative analysis experiment. To assess the retention of geometric attributes and semantic features from structure images in generated images, we employ three key indicators: Structure Preservation (SP), Chamfer Distance (CD), and Fréchet Inception Distance (FID). Concurrently, to evaluate the fidelity of low-level features between appearance images and generated images, we utilize three distinct metrics: Gray-Level Co-occurrence Matrix (GLCM), Peak Signal-to-Noise Ratio (PSNR), and Color Histogram Correlation (CHC).
  • Structure Preservation (SP): we utilize the marquee interaction mode of SAM [23] for selecting areas to obtain binary masks corresponding to structure images and their respective output images. Then, we compute their Intersection over Union (IoU) results as a measure of Structure Preservation.
  • Chamfer Distance (CD): we first extract the line drawings of the structure and generated images, and then filter out redundant details using the Canny operator. The high and low thresholds used by the Canny operator are set to 150 and 50, respectively. Finally, we calculate the chamfer distance between the line drawings as a measure of the gap between the sets of edge points in the two images. A smaller value indicates a higher degree of match between the shape or edge features in the structure and generated images.
  • Fréchet Inception Distance (FID): we calculate the FID score between the structure image and the generated image to quantify the extent to which the two images align in terms of their structural features.
  • Gray-Level Co-occurrence Matrix (GLCM): it is used to calculate the loss value of texture features between the appearance image and the generated image.
  • Peak Signal-to-Noise Ratio (PSNR): it is used to measure how well the generated image preserves the low-level features of appearance.
  • Color Histogram Correlation (CHC): it is used to calculate the color similarity between the generated image and the appearance image. Among them, we use mask to cover the background of the image.
Test Datasets: AFHQ [10], CUB-200-2011 [44], FGVC-Airaft [32], Stanford-Cars [24]. Among them, the three domain data of cat, dog, and wild are all from the AFHQ test set. We followed the setting of AFHQ’s test set, with 500 images for each category, and randomly selected 500 images from three other datasets as appearance images. For every types of appearance transfer tasks there are 2000 pairs, Therefore, the data in Table 5 is based on the evaluation results of 12000 Structure-Appearance image pairs.
5.2 Qualitative analysis
The qualitative analysis experiments include a total of eight domains (cat, dog, wild, bird, airplane, car, bronze, architecture). In addition to the datasets used in the quantitative analysis experiments, additional datasets are: WikiChurches [3] and Bronze Dings [54].
Among them, we mainly designed two types of tasks. One is the style transfer as shown in Figure 10 including three domains (bronze, airplane, architecture). In this task, the style image and the content image do not have semantic correspondence, and their relationships belong to different domains. So the focus in this task is to obtain the low-level texture features from the style image without semantic correspondence and then render it to the content image.
Figure 10: Style transfer tasks. Our focus is to obtain the low-level texture features from the style image without semantic correspondence and then render it to the content image.
The another is the appearance transfer task as shown in Figure 11 that includes six domains (cat, dog, wild, bird, airplane, car). The image pairs used in this task belong to the same data distribution, and focus is to transfer the underlying color distribution and stroke texture in the appearance image to the structure image without destroying its geometric properties.
Figure 11: Appearance transformation tasks. A total of six domains (cat, dog, wild, bird, airplane, car), we showcase the effects achieved by the baseline architecture with filtering combined operation in local, object-centric, and full-image-level appearance transfer tasks.
6 Limitation and Conclusion
While the field of controllable generation tasks is flourishing, feature decomposition remains a formidable challenge. Therefore, harnessing distinctions in distribution and frequency domains among features in pixel space, we introduce FilterPrompt, an approach designed to amplify the efficacy of model control. This approach boasts universal applicability across any diffusion model, enabling users to adjust the representation of specific image features to task specifications, thereby facilitating more precise and controllable generation outcomes. Through our experiments, we demonstrate that utilizing combinations such as the ITV-R 601-2 luma transform for decolorization, inversion, and contrast enhancement of grayscale images can heighten the diffusion model’s control over structural features, while Gaussian FilterPrompt can mitigate the expression of redundant texture features.
Beyond the baseline framework utilized in this study, the integration of other advanced diffusion models may potentially lead to even superior results if they are implemented to replace certain components in the baseline framework. However, exploring this possibility is beyond the scope of our current discussion. Notably, the design of FilterPrompt requires manual intervention in accordance with the task requirements and data characteristics. This process entails making preliminary judgments regarding the focus of content information and incurring certain trial-and-error costs.
In summary, we contend that the integration of FilterPrompt with diffusion models can achieve remarkably impressive control effects compared to alternative approaches within the feature space. Moreover, it is exceedingly convenient, making it suitable for widespread adoption in other diffusion models.